OpenStudy (anonymous):

Select the product of (-4x7y)(x3y4). confused how to do these

OpenStudy (shamil98):

\[(-4x^7y)(x^3y^4)\] Multiply the terms... remember the rule for multiplying exponents: \[\large x^n * x^m = x^{n+m}\]

OpenStudy (shamil98):

\[-4x^7 * x^3 * y*y^4 = ?\]

OpenStudy (anonymous):

-4x^7 *x^3 * y^5? derp

OpenStudy (shamil98):

yes, keep going..

OpenStudy (anonymous):

hmm is it, -4x^10y^5 ??

OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

oh haha thanks!

OpenStudy (anonymous):

hmm so how about this one? Select the product of (9x + 6)(9x - 6).

OpenStudy (shamil98):

multiply them using the FOIL method.

OpenStudy (anonymous):

i just learned about the FOIL jutsu, can u explain to me how to do it :/

OpenStudy (shamil98):

or you could remember the identity: (x+a)(x-a) = x^2 - a^2

OpenStudy (shamil98):

(9x+6)(9x-6) = 81x^2 - 36

OpenStudy (anonymous):

oh okay, the exponent always throws me off

OpenStudy (anonymous):

but what do you do when it gets more complicated like this Select the product of (6x + 9)(x2 - 4x + 4).

OpenStudy (anonymous):

i lose count of all the crap in there

OpenStudy (shamil98):

all you do is multiply each term to the terms in the parentheses.. (6x+9)(x^2 -4x +4) = 6x*x^2 + 6x*(-4x) + 6x*4 + 9*x^2 + 9*(-4x) + 9*4 = .....

OpenStudy (anonymous):

6x^3-15^2-12x+36

OpenStudy (shamil98):

no.

OpenStudy (anonymous):

fudge

OpenStudy (anonymous):

is this some sort of genjutsu?

OpenStudy (shamil98):

6x*x^2 + 6x*(-4x) + 6x*4 + 9*x^2 + 9*(-4x) + 9*4 = ..... 6x^3 - 24x^2 + 24x + 9x^2 - 36x + 36 = 6x^3 - 15x^2 - 12x + 36

OpenStudy (shamil98):

you forgot the x in -15x^2 bud

OpenStudy (anonymous):

oh lol -.-

OpenStudy (anonymous):

then i was right lmao u cofused me